STRUCTURES AND SPECTROSCOPIC PROPERTIES CALCULATED FOR C_{6}H_{7}^{+} AND ITS COMPLEXES WITH Ne, Ar, N_{2}, OR CO_{2}

\underline{P. BOTSCHWINA} and R. OSWALD, Institute of Physical Chemistry, University of Göttingen, Tammannstr. 6, D-37077 Göttingen, Germany.

Explicitly correlated coupled cluster theory at the CCSD(T)-F12x (x = a, b) level^{a} in conjunction with the double-hybrid density functional B2PLYP-D^{b} has been employed in a study of the benzenium ion (C_{6}H_{7}^{+}) and its complexes with simple ligands (L = Ne, Ar, N_{2}, or CO_{2}).^{c} The ground-state rotational constants of C_{6}H_{7}^{+} are predicted to be A_{0} = 5445 MHz, B_{0} = 5313 MHz, and C_{0} = 2731 MHz. For the complexes with L = Ne, Ar or N_{2}, the energetically most favourable structure is of π-bonded type, but for the most strongly bound complex C_{6}H_{7}^{+} · CO_{2} a conformer with the CO_{2} ligand lying in the ring-plane of the C_{6}H_{7}^{+} moiety is slightly lower in energy.

aT. B. Adler, G. Knizia, and H.-J. Werner, J. Chem. Phys. \underline{127}, 221106 (2007);

G. Knizia, T. B. Adler, and H.-J. Werner, J. Chem. Phys. \underline{130}, 054104 (2009).

bT. Schwabe and S. Grimme, Phys. Chem. Chem. Phys. \underline{9}, 3397 (2007).

cP. Botschwina and R. Oswald, J. Phys. Chem. A \underline{115}, 13664 (2011);

P. Botschwina and R. Oswald, J. Chem. Phys. submitted.